A train approaching a platform at a speed of 54 km h-1 sounds a whistle. An observer on the platform finds its frequency to be 1620 Hz. The train passes the platform keeping the whistle on and without slowing down. What frequency will the observer hear after the train has crossed the platform? The speed of sound in air = 332 m s-1.

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. Let the frequency of the bullet will be f Given, u = 330 m/s, v base s = 220 m/s a) Apparent frequency before crossing = f’ = (330/330 - 220) f = 3f b) Apparent frequency after crossing = f” = (330/530 - 220) f = 0.6f So, (f”/f’) = 0.6f/3f = 0.2 Therefore, fraction change = 1 – 0.2 = 0.8